Hi
layla spent a total of $135 on a skirt and shoes.the price of the shoes was $18 less than twice what she paid for the skirt.find the cost of the shoes.
thank you

Answer by Cromlix(4381)

Hi there,
Make the skirt = x
Shoes = 2x - 18
x + 2x - 18 = 135
3x - 18 = 135
3x = 135 + 18
3x = 153
x = 153/3
x = $51
Skirt cost $51.00
Shoes cost $84.00
Hope this helps :-)